# Certified Randomness implies Secure Classical Position-Verification

### v2

@@ 1.2 Paper Organization
In Section 1.3 we outline the main technical ideas. We recall some preliminary concepts and definitions useful for relevant to our in Section 2. In Section 3, we provide our compiler for CVPV from single-round certified randomness and prove the security of the construction, our first major result. In Section 4 we provide a number of methods of generalizing this compiler to allow for multi-round certified randomness protocols. Specifically, in Section 4.1 we show that certified randomness protocols with a property we refer to as sequential decomposability suffice to allow for a natural generalization of our compiler for multi-round settings, and in Section 4.2 we prove (with a full proof in Appendix A) that that a class of natural certified randomness protocols satisfy the necessary property. In Section 4.3, and Section 4.4 we prove the security of an alternate compilation method, based on additional timing constraints, which has advantages in the idealized model at the cost of practical robustness. Finally, in Appendix B, we show that the well known near-term proposal for certified randomness due to Aaronson and Hung suffices to instantiate our compiler.
1.3 Technical Overview
Single Round Compiler
We first consider any single-round PoQ-based certified randomness protocol $\mathscr{P}=(P,V)$, which is modeled as a two-message interactive protocol. In the first message, the verifier sends a random challenge $\mathsf{ch}$ to the prover. In the second message, it gets back $\mathsf{ans}$ as a response. Later, the verifier performs a verification process $\mathsf{Ver}$ to accept or reject the response.
Figure 1: A spacetime diagram describing the honest interaction between two verifiers at locations $X=0$ and $X=2$ and an honest prover located at the claimed position (for simplicity, located at $X=1$ in the diagram). The verifiers pre-share random values $x,y,k,r$ prior to the start of the protocol, and set $s=ch\oplus G_{k}(x\oplus y)$ for some family of hash functions $\{G_{k}\}$ and $ch=\mathsf{Gen}(1^{n};r)$ for $\mathsf{Gen}$ the verifier functionality of some certified randomness verifier and where $accept(\cdot)$ determines a boolean function that determines whether the verifiers accept the provers claim. In all spacetime figures we provide, we implicitly require that the proper timing constraints are observed in order for the verifiers to accept.
Figure 2: A spacetime diagram describing the honest interaction between two verifiers at locations $X=0$ and $X=2$ and two dishonest prover located at arbitrary positions between $V_{1}$ and $V_{2}$ (for simplicity, located at $X=0.5$ and $X=1.5$ in the diagram). The provers are allowed to communicate during the protocol but are non-signaling. Notation and behavior is otherwise analogous to Figure 1. In particular, verifiers expect to recieve messages at times determined by honest prover’s position.
Construction: In our compiler, we consider that the two verifiers (say $V_{1},V_{2}$) in the CVPV protocol have access to a family of cryptographic hash function $\left\{G_{k}\right\}_{k\in\{0,1\}^{\lambda}}:\{0,1\}^{m}\to\{0,1\}^{n}$. In addition, they share a random hash key $k$ and two random inputs $x,y\leftarrow\{0,1\}^{m}$. At time $t=0$, $V_{1}$ sends $s:=G_{k}(x\oplus y)\oplus\mathsf{ch}$, $x,k$ to the prover $P$ at a claimed location (say $X=1$ in Figure 1) and $V_{2}$ sends $y$ to $P$. Upon receiving $x,s,k,y$ from the verifiers, the prover can recover $\mathsf{ch}$ and send back the response $\mathsf{ans}$ to the verifiers. The verifiers verify the position of the prover if the following two conditions are satisfied.
  1. 1.
    Timing Constraint: Both verifiers receive at time $t=2$.
  2. 2.
    Consistency Constraint: Both of the received answers are the same.
  3. 3.
    Certified Randomness Constraint: The verification process $\mathsf{Ver}$ corresponding to the certified randomness protocol accepting the response.
Soundness Proof Sketch (Proof Sketch of Theorem 1).
@@ 1.3 Technical Overview
We reduce the soundness of the CVPV protocol above to the certified randomenss property of $\mathscr{P}$. We first make the connection to certified randomness by imagining an ideal game, where $P_{1}$ and $P_{2}$ are simultaneously given $\mathsf{ch}$ and asked to output the same correct answer $\mathsf{ans}$. If no communication is allowed, then beating this game violates certified randomness due to no-signalling. Intuitively, certified randomness dictates that $\mathsf{ans}$ cannot be computed deterministically from $\mathsf{ch}$, which forces $P_{1}$ and $P_{2}$ to perform their own local random computation to output a valid answer.
@@
One would hope that this in turn hides any useful information about $\mathsf{ch}$ during the same timeframe. Nonetheless, it is not clear how to argue this directly. Namely, one needs to rule out homomorphic attacks, where $P_{1}$ performs a quantum computation on input $x$, $P_{2}$ on input $y$, and then they can each deterministically recover the same output $\mathsf{ans}$, which could be obtained by running the honest prover of $\mathscr{P}$ on input $\mathsf{ch}$.
To circumvent this issue, we use a cryptographic hash function $G$ and encrypt $\mathsf{ch}$ with a one-time-pad using $G(x\oplus y)$. The property we need from $G$ is query-extractability, and accordingly we show security in the quantum random oracle model.
Multi-Round Sequential Compiler
@@
Construction: The compiler corresponding to such an $\ell$-round certified randomness protocol is a sequential repetition of the interactive portion of the single-round compiler, followed by the necessary testing of the entire transcript. Before the beginning of the protocol, i.e., at $t=-\infty$ the verifiers ($V_{1},V_{2}$) share $\ell$ random hash keys $\{k_{i}\}_{i\in[\ell]}$, $\ell$ random input pairs $\{(x_{i},y_{i})\}_{i\in[\ell]}$, and $\ell$ random challenges $\{\mathsf{ch}_{i}\}_{i\in[\ell]}$ corresponding to the challenges of the certified randomness protocol. On the $i\in[\ell]$ round, $V_{1}$ sends $s_{i}:=G_{k_{i}}(x_{i}\oplus y_{i})\oplus\mathsf{ch}_{i}$ and $V_{2}$ sends $y_{i}$ to the prover $P$ at a claimed location $X=1$ (see Figure 3 for reference) at time $t_{i-1}$. Upon receiving $k_{i},x_{i},s_{i},y_{i}$, the prover computes $\mathsf{ch}_{i}$, and sends back the answer $\mathsf{ans}_{i}$ corresponding to the challenge $\mathsf{ch}_{i}$ to the verifiers. Suppose, the verifiers send the challenges at time $t^{\text{send}}_{i}$, and receive the answers at time $t^{\text{rec}}_{i}$. At the end of the protocol, the verifiers accepts the claimed location of the prover if the answers satisfy the timing constraint, i.e. $t^{\text{rec}}_{i}-t^{\text{send}}_{i}=2$ for all $i\in[\ell]$, the consistency check, i.e. $\mathsf{ans}_{i}=\mathsf{ans}^{\prime}_{i}$ for all $i\in[\ell]$, and the certified randomness constraint. We refer to Figure 3 for the schematic diagram of the protocol and Figure 4 for the behavior of cheating provers.
@@
Difficulty of the Soundness Proof. The adversarial model of all the existing multi-round certified randomness protocols [AH23, BCM+21a] use entropy accumulation theorem (EAT) to calculate a lower bound on the min-entropy of the produced outcomes for all the rounds. More precisely, the multi-round certified randomness protocol in [AH23] models the entire adversary channel as an entropy accumulation channel (EAT) [DFR20]. Although the original security analysis in [BCM+21a] did not directly use EAT, later, Merkulov and Arnon-Friedman show in [MAF23] that the adversarial channel of [BCM+21a] can indeed be modeled as an EAT channel. The EAT channel that is proposed in [DFR20] does not allow the prover to communicate its private registers to the external adversary during the runtime of the protocol. However, in our sequential compiler, after the $i$-th round the malicious provers can communicate with each other, and exchange their internal registers as well as their answers for the $i$-th round. We refer to Figure 4 for an example. This stops us from applying the security analysis from [DFR20, BCM+21a, AH23] directly. As a way out, we require that our certified randomness protocols satisfy a stronger notion of security, namely sequential decomposition that we define informally in Definition 1 (formally stated in Definition 5).
Soundness Proof Sketch (Proof Sketch of Theorem 2).
Similar to the soundness proof of the single round protocol, here we also reduce the soundness of the multi-round CVPV protocol to the multi-round certified randomness protocol. Here, the only difference is that due to the multi-round nature of the compiler, we need to consider a multi-round ideal guessing game. In this ideal guessing game on the $i$-th round $P_{1}$ and $P_{2}$ are simultaneously given $\mathsf{ch}_{i}$ and asked to output the same correct answer $\mathsf{ans}_{i}$. Note that, the usual definition of the certified randomness do not provide any guarantee on the winning probability of this guessing game. Therefore, we need to add the sequential decomposition property. Indeed, if no communication is allowed then due to the sequential decomposition property of the multi-round certified randomness protocol the winning probability of this guessing game will be negligible. The rest of the reduction to the real security game is similar to the proof sketch of Theorem 1. We refer to Theorem 7 for a more detailed analysis.
Sequential Decomposition from Repetition
@@
We strengthen our result on multi-round compilers by showing the existence of a family of certified randomness protocols that are secure under our notion of sequential decomposition (Definition 1 or Definition 5 for a more formal version). In Theorem 3, we prove that certified randomness protocols based on repetition with single round entropy guarantees, including the well known protocols in [AH23, BCM+21a], satisfy Definition 5.
@@
Proof Sketch of Theorem 3 We prove Theorem 3 by using the key observation that the prover’s answers from the previous rounds do not help the guesser after the guess is already committed for each round. First, we consider an optimal adversary (maximum probability of succeeding the proof of quantumness test of the certified randomness protocol and the consistency check of the sequential decomposition property) where the prover may send information about previous answers to the guesser. We then consider a slightly modified adversary that has the same optimal success probability, but only the guesser is allowed to send information to the prover. For this modified adversary, the guesser simply assumes that all guesses are correct thus far. If at least one guess is incorrect, the protocol has already failed and the adversary strategy from then on does not matter. If the all guesses are correct, then pretending the answers are always the same as the guesses results in the correct behavior.
@@
Rapid Fire Compiler
@@
Soundness Proof Sketch. For the rapid-fire compiler, if we assume that the malicious provers do not communicate the answers to each other during the runtime of the protocol, then we can directly apply the entropy guarantee to upper bound the guessing probability without the sequential decomposition property. However, to satisfy this requirement we need to assume the condition $\Delta l<2t_{\text{comm}}$ applies to the protocol, where $t_{\text{comm}}$ denotes the communication time between any two malicious provers.
@@
Instantiation using Random Circuit Sampling (RCS)
We instantiate the protocol using certified randomness based on random circuit sampling (RCS). Aaronson and Hung [AH23] show that under some models (e.g. a fully general adversary given oracle access to challenge circuits), passing the XHOG test (achieving high average probability $p_{C}(z)$ over choices of $C$ and output $z$) implies $\Omega(n)$ von Neumann entropy conditioned on quantum side information, where $n$ is the number of qubits. We formally define the protocols in Appendix B, shown in Figure 9 and Figure 10. The single round entropy guarantees are given by Theorem 12 and Theorem 13.
1.4 Open Problems
In this section, we highlight the remaining problems related to classical verifier position verification.
Security in the Plain Model:
In this paper, we prove the soundness of the compilers in quantum random oracle model. However, it is crucial to have the security analysis in the plain model for a practical implementation of this protocol.
Minimum Sufficient Condition for the CVPV:
This paper requires proof of quantumness-based certified randomness to construct the CVPV protocol. However, it is not clear whether this is the most minimal assumption. We leave the minimum sufficient assumption required to construct a CVPV protocol as an open problem.
@@ 2 Preliminaries
We now recall a collection of useful definitions and results froim the literature. Throughout the paper, we denote the security parameter by $\lambda$.
@@ 2.1 Quantum Information
2.2 Proof of Quantumness (PoQ)
@@
  • $s$-Soundness: For any PPT prover $P$, the probability that $V$ accepts is at most $s+\mathsf{negl}(\lambda$).
Certified Randomness.
A property stronger than soundness is certified randomness. At a high level, certified randomness requires that if a prover $P$ passes the PoQ protocol, then its output cannot be predicted by a guesser $E$ who:
@@
    Shares entanglement with $P$.
In the most general case, we require this to hold even if $P$ is fully malicious (a.k.a. a fully general device). We give the formal definition below.
Definition 3 (Certified Randomness).
A PoQ protocol $\mathscr{P}$ is said to have certified randomness property if for any QPT prover $P$, conditioned on $V$ accepting, it is true with overwhelming probability that
$H_{\min}({\color[rgb]{1,.5,0}Z}\;|\;{\color[rgb]{1,.5,0}CE})_{\rho}\geq\omega\left(\log\lambda\right),~{}~{}\text{where,}$,
  • $\rho=\rho_{{\color[rgb]{1,.5,0}ZCDE}}$ is the final state of $P$, with ${\color[rgb]{1,.5,0}D}$ being the working register.
  • ${\color[rgb]{1,.5,0}C}$ is a classical register containing all the messages sent by $V$.
  • ${\color[rgb]{1,.5,0}A}$ is the classical register containing the final message sent by $P$.
  • ${\color[rgb]{1,.5,0}E}$ is arbitrary (possibly entangled) quantum side information of $P$ such that $P$ does not act on ${\color[rgb]{1,.5,0}E}$ during the protocol.
@@ 2.3 Position Verification with Classical Verifiers
A position verification scheme $\mathscr{V}=(V,X)$ with classical verifiers is an interactive relativistic protocol between a set of classical verifiers $V=(V_{i},X_{i})_{i\in I}$, where each classical verifier $V_{i}$ is located at position $X_{i}\in\mathbb{R}$ on the real line, and $X\in\mathbb{R}$ is the purported location of a quantum prover. During the protocol, we assume that all communication happens at the speed of light (1) and all computation is instantaneous. We also assume that $V$ can perform a secure setup before the protocol, and communicate securely during the protocol.

### v3

@@ 1.6 Paper Organization
We give an overview of our techniques in Section 2. We recall some relevant preliminary concepts in Section 3. In Section 4, we provide our compiler for CVPV from single-round certified randomness and prove the security of the construction, our first major result. In Section 5 we provide a number of methods of generalizing this compiler to allow for multi-round certified randomness protocols. Specifically, in Section 5.2 we show that certified randomness protocols with a property we refer to as sequential decomposability suffice to allow for a natural generalization of our compiler for multi-round settings, and in Section 5.3 we prove (with a full proof in Section 7) that a class of natural certified randomness protocols satisfy the necessary property. In Section 5.4, and Section 5.5 we prove the security of an alternate compilation method, based on additional timing constraints, which has advantages in the idealized model at the cost of practical robustness. In Section 8.1, we show that the well known near-term proposal for certified randomness due to Aaronson and Hung suffices to instantiate our compiler. The rest of Section 8 shows other instantiations of our compiler that are not NISQ-friendly but theoretically relevant.
Acknowledgments
We thank Dakshita Khurana and Kabir Tomer for pointing out an inaccuracy in our citation of recent work. We thank James Bartusek for valuable discussions regarding the implications of our results.
Disclaimer
This paper was prepared for informational purposes with contributions from the Global Technology Applied Research center of JPMorgan Chase & Co. This paper is not a product of the Research Department of JPMorgan Chase & Co. or its affiliates. Neither JPMorgan Chase & Co. nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
2 Technical Overview
2.1 Single Round Compiler
The informal construction and schematics of the single round compiler can be found in the main text, so we omit them here.
@@ 2.1 Single Round Compiler
Soundness Proof Sketch (Proof Sketch of Theorem 1). We reduce the soundness of the CVPV protocol above to the certified randomenss property of $\mathscr{P}$. We first make the connection to certified randomness by imagining an ideal game, where $P_{1}$ and $P_{2}$ are simultaneously given $\mathsf{ch}$ and asked to output the same correct answer $\mathsf{ans}$. If no communication is allowed, then beating this game violates certified randomness due to no-signalling. Intuitively, certified randomness dictates that $\mathsf{ans}$ cannot be computed deterministically from $\mathsf{ch}$, which forces $P_{1}$ and $P_{2}$ to perform their own local random computation to output a valid answer.
@@
One would hope that this in turn hides any useful information about $\mathsf{ch}$ during the same timeframe. Nonetheless, it is not clear how to argue this directly. Namely, one needs to rule out homomorphic attacks, where $P_{1}$ performs a quantum computation on input $x$, $P_{2}$ on input $y$, and then they can each deterministically recover the same output $\mathsf{ans}$, which could be obtained by running the honest prover of $\mathscr{P}$ on input $\mathsf{ch}$. This issue was previously faced by Unruh [Unr14] in the context of quantum position verification, where the classical basis information had to be hidden from the adversary in this strong sense.
To circumvent this issue, we follow [Unr14] and use a cryptographic hash function $G$ and encrypt $\mathsf{ch}$ with a one-time-pad using $G(x\oplus y)$. The property we need from $G$ is query-extractability, and accordingly we show security in the quantum random oracle model. This technique was also used by Liu et al. [LLQ22] to achieve security against unbounded entanglement.
2.2 Multi-Round Sequential Compiler
@@
Construction: The compiler corresponding to such an $\ell$-round certified randomness protocol is a sequential repetition of the interactive portion of the single-round compiler, followed by the necessary testing of the entire transcript. Before the beginning of the protocol, i.e., at $t=-\infty$ the verifiers ($V_{1},V_{2}$) share $\ell$ random hash keys $\{k_{i}\}_{i\in[\ell]}$, $\ell$ random input pairs $\{(x_{i},y_{i})\}_{i\in[\ell]}$, and $\ell$ random challenges $\{\mathsf{ch}_{i}\}_{i\in[\ell]}$ corresponding to the challenges of the certified randomness protocol. On the $i$-th ($i\in[\ell]$) round, $V_{1}$ sends $s_{i}:=G_{k_{i}}(x_{i}\oplus y_{i})\oplus\mathsf{ch}_{i}$ and $V_{2}$ sends $y_{i}$ to the prover $P$ at a claimed location $X=1$ (see Figure 2 for reference). Upon receiving $k_{i},x_{i},s_{i},y_{i}$, the prover computes $\mathsf{ch}_{i}$, and sends back the answer $\mathsf{ans}_{i}$ corresponding to the challenge $\mathsf{ch}_{i}$ to the verifiers. Suppose, the verifiers send the challenges at time $t^{\text{send}}_{i}$, and receive the answers at time $t^{\text{rec}}_{i}$. At the end of the protocol, the verifiers accept the claimed location of the prover if the answers satisfy the timing constraint, i.e. $t^{\text{rec}}_{i}-t^{\text{send}}_{i}=2$ for all $i\in[\ell]$, the consistency check, i.e. $\mathsf{ans}_{i}=\mathsf{ans}^{\prime}_{i}$ for all $i\in[\ell]$, and the certified randomness constraint. We refer to Figure 2 for the schematic diagram of the protocol and Figure 3 for the behavior of cheating provers.
@@
Difficulty of the Soundness Proof. The adversarial model of all the existing multi-round certified randomness protocols [AH23, BCM+21] use entropy accumulation theorem (EAT) to calculate a lower bound on the min-entropy of the produced outcomes for all the rounds. More precisely, the multi-round certified randomness protocol in [AH23] models the entire adversary channel as an entropy accumulation channel (EAT) [DFR20]. Although the original security analysis in [BCM+21] did not directly use EAT, later, Merkulov and Arnon-Friedman show in [MAF23] that the adversarial channel of [BCM+21] can indeed be modeled as an EAT channel. The EAT channel that is proposed in [DFR20] does not allow the prover to communicate its private registers to the external adversary during the runtime of the protocol. However, in our sequential compiler, after the $i$-th round the malicious provers can communicate with each other, and exchange their internal registers as well as their answers for the $i$-th round. We refer to Figure 3 for an example. This stops us from applying the security analysis from [DFR20, BCM+21, AH23] directly. As a way out, we require that our certified randomness protocols satisfy a stronger notion of security, namely sequential decomposition that we define informally in Definition 1 (formally stated in Definition 5). Soundness Proof Sketch (Theorem 2). Similar to the soundness proof of the single round protocol, here we also reduce the soundness of the multi-round CVPV protocol to the multi-round certified randomness protocol. Here, the only difference is that due to the multi-round nature of the compiler, we need to consider a multi-round ideal guessing game. In this ideal guessing game on the $i$-th round $P_{1}$ and $P_{2}$ are simultaneously given $\mathsf{ch}_{i}$ and asked to output the same correct answer $\mathsf{ans}_{i}$. Note that, the usual definition of the certified randomness do not provide any guarantee on the winning probability of this guessing game. Therefore, we need to add the sequential decomposition property. Indeed, if no communication is allowed then due to the sequential decomposition property of the multi-round certified randomness protocol the winning probability of this guessing game will be negligible. The rest of the reduction to the real security game is similar to the proof sketch of Theorem 1. We refer to Theorem 9 for a more detailed analysis.
2.3 Sequential Decomposition from Repetition
@@
We strengthen our result on multi-round compilers by showing the existence of a family of certified randomness protocols that are secure under our notion of sequential decomposition (Definition 1 or Definition 5 for a more formal version). In Theorem 3, we prove that certified randomness protocols based on repetition with single round entropy guarantees, including the well known protocols in [AH23, BCM+21], satisfy Definition 5.
@@
Proof Sketch (Theorem 3). We prove Theorem 3 by using the key observation that the prover’s answers from the previous rounds do not help the guesser after the guess is already committed for each round. First, we consider an optimal adversary (maximum probability of succeeding the proof of quantumness test of the certified randomness protocol and the consistency check of the sequential decomposition property) where the prover may send information about previous answers to the guesser. We then consider a slightly modified adversary that has the same optimal success probability, but only the guesser is allowed to send information to the prover. For this modified adversary, the guesser simply assumes that all guesses are correct thus far. If at least one guess is incorrect, the protocol has already failed and the adversary strategy from then on does not matter. If the all guesses are correct, then pretending the answers are always the same as the guesses results in the correct behavior.
@@
2.4 Rapid Fire Compiler
@@
Soundness Proof Sketch (Theorem 4). For the rapid-fire compiler, if we assume that the malicious provers do not communicate the answers to each other during the runtime of the protocol, then we can directly apply the entropy guarantee to upper bound the guessing probability without the sequential decomposition property. However, to satisfy this requirement we need to assume the condition $\Delta l<2t_{\text{comm}}$ applies to the protocol, where $t_{\text{comm}}$ denotes the communication time between any two malicious provers.
@@ 3 Preliminaries
We now recall a collection of useful definitions and results from the literature. Throughout the paper, we denote the security parameter by $\lambda$.
@@ 3.1 Quantum Information
3.2 Proof of Quantumness (PoQ) and Certified Randomness
@@
  • Soundness: For any PPT prover $P$, the probability that $V$ accepts is at most $\mathsf{negl}(\lambda$).
A single-round PoQ protocol is a two-message PoQ protocol.
Certified Randomness. A property stronger than soundness is certified randomness. At a high level, certified randomness requires that if a prover $P$ passes the PoQ protocol, then its output cannot be predicted by a guesser $E$ who:
@@
    Shares entanglement with $P$, but does not communicate with $P$ during the protocol.
In the most general case, we require this to hold even if $P$ is fully malicious (a.k.a. a fully general device). We give the formal single-round definition below. The definitions for the multi-round case are given in Section 5.
Definition 3 (Certified Randomness (Single-Round)).
A PoQ protocol $\mathscr{P}$ is said to have certified randomness property if no pair of a QPT prover $P$ and a (possibly entangled) unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
  1. 1.
    The verifier $V$ of $\mathscr{P}$ sends a challenge $\mathsf{ch}$ to both $P$ and $Q$.
  2. 2.
    $P$ sends back an answer $\mathsf{ans}$ and $Q$ and outputs a guess $\mathsf{ans}^{\prime}$.
  3. 3.
    $(P,Q)$ win if $V$ accepts and $\mathsf{ans}=\mathsf{ans}^{\prime}$.
Remark 1.
Definition 3 captures certified randomness generation using a single device ($P$). The guesser $Q$, who is not allowed to communicate with $P$, appears in the definition simply because it is a game-based definition that is based on the operational meaning of min-entropy. Intuitively, the guarantee is that $P$ generates $\mathsf{ans}$ with high min-entropy conditioned on any quantum side information held by $Q$. This is in contrast with certified randomness generation based on Bell inequality violations [PAM+10], which require two devices and the no-communication assumption during the real execution of the protocol.
@@ 3.3 Position Verification with Classical Verifiers
A position verification scheme $\mathscr{V}=(V,X)$ with classical verifiers is an interactive relativistic protocol between a set of classical verifiers $V=(V_{i},X_{i})_{i\in I}$, where each classical verifier $V_{i}$ is located at position $X_{i}\in\mathbb{R}$ on the real line, and $X\in\mathbb{R}$ is the purported location of a quantum prover. During the protocol, we assume that all communication happens at the speed of light (set to be $1$) and all computation is instantaneous. We also assume that $V$ can perform a secure setup before the protocol, and communicate securely during the protocol.
@@
Remark 2 (Efficient Verification).
Note that although it is a desirable property, we do not require the verifier(s) to be efficient in Definitions 2, 3 and 4. The reason is that our NISQ-friendly instantiation of CVPV in Section 8.1 is based on RCS which does not have efficient verification. The rest of the protocols we consider in this work have efficient verification.